OpenStudy (anonymous):

Find an equation for the line that passes through the points (3,2) and (-3,5) .

OpenStudy (anonymous):

use the errr.I forgot the name of the form. But its is: Y-y1=m(x-x1)

OpenStudy (cathyangs):

that's point slope. the slope would be m, and in this case, (2-5)/(3-(-3)), this simplifies to -3/6, or -1/2. use this as m and take the x and y of one of those points, and use them as y1 and x1 in jhieuluong's formula. You can do some algebra, and get to slope-intercept, like y=mx+b form.